Hi, This question is regarding Supply chain management.

Amaro Pte Ltd, a battery operated vacuum for consumer market has a problem. its customers, Wallmart and Goldspice both wants exclusive products, so that consumers cannot make price comparisons of their products. For example, Amaro pte ltd sells similar car-vacuum to each customer buts has to use packaging customized to each retailer(including a different model number). Supposing the weekly demand for the vacuum cleaner to each retailer is normal with mean 5,000 & standard deviation of 2,200. Amaro Pte Ltd cost of doing set-up to produce a new lot is $5,000 & cost of holding a vacuum cleaner is stock for a year is $52.00. Each lot takes a week to produce. As the two retailers are quite important to Amaro Pte Ltd, it sets a target fill rate of 99.7%

Quetion:

What is Amaro Pte Ltd Average inventory of each of the two versions of this vacuum cleaner, assuming the company follows (R,Q) policy?
Answer #1

Weekly demand, d = 5000

Annual demand, D = 5000*52 = 260,000

Standard deviation of weekly demand, s = 2200

Setup cost, K = $ 5000

Holding cost, H = $ 52

Lead time, L = 1 week

Target fill rate, = 99.7 %

EOQ = sqrt(2DK/H)

= sqrt(2*260000*5000/52)

= 7,071 units

L(z) = (1-)*Q/s = (1-0.997)*7071/2200 = 0.0096

Corresponding value of z = 1.955

Safety stock, ss = z*s = 1.955*2200 = 2,074 units

Reorder point, R = d*L + ss = 5000*1 + 2074 = 7,074 units

Average inventory = Q/2 + ss = 7071/2 + 2074 = 5,610 units